# Quantum HodgeRank: Topology-Based Rank Aggregation on Quantum Computers

### v2

@@ Abstract
where $w_{ij}$ represents the weight associated to the data on (oriented) edge $[v_{i},v_{j}]$ and $\bm{\mathrm{B}}_{k}$ is a matrix representation of the $k$-th boundary operator of a clique complex $\mathcal{K}_{n}(\mathcal{G})$ (see Appendix I). The space $\mathrm{im}(\bm{\mathrm{B}}_{1}^{\dagger})$ contains so-called conservative edge flows, which means there exists $\bm{\mathrm{s}}^{0}_{*}\in\mathbb{R}^{|\mathcal{V}|}$ such that $\bm{\mathrm{s}}^{1}_{\mathrm{G}}([v_{i},v_{j}])=\bm{\mathrm{s}}^{0}_{*}(v_{i})-\bm{\mathrm{s}}^{0}_{*}(v_{j})$. In a discrete exterior calculus sense, $\bm{\mathrm{s}}^{1}_{\mathrm{G}}$ is an analogue of the gradient of the potential function $\bm{\mathrm{s}}^{0}_{*}$.
HodgeRank then outputs the potential (or score) function $\bm{\mathrm{s}}^{0}_{*}$ from Eq. (1) by $\bm{\mathrm{B}}_{1}^{\dagger}\bm{\mathrm{s}}^{0}_{\ast}=\bm{\mathrm{s}}^{1}_{\mathrm{G}}$. This defines a global ranking of the alternatives: the score of alternative $v_{i}$ is $\bm{\mathrm{s}}^{0}_{*}(v_{i})$, and $v_{i}$ outranks $v_{j}$ whenever $\bm{\mathrm{s}}^{0}_{*}(v_{i})\geq\bm{\mathrm{s}}^{0}_{*}(v_{j})$. In this Letter, we set $w_{ij}=0$ if $v_{i},v_{j}$ are not compared, and 1 otherwise; we discuss this assumption in Appendix III. Under this assumption, $\bm{\mathrm{s}}^{0}_{\ast}=\big{(}\bm{\mathrm{B}}_{1}\bm{\mathrm{B}}_{1}^{\dagger}\big{)}^{+}\bm{\mathrm{B}}_{1}\bm{\mathrm{s}}^{1}$, where $+$ denotes the Moore-Penrose pseudoinverse. The relative length of $\bm{\mathrm{s}}^{1}_{\mathrm{G}}$ compared to $\bm{\mathrm{s}}^{1}$ measure the rankability of the alternatives according to $\bm{\mathrm{s}}^{1}$.
@@
These algorithms do not lead to efficient quantum algorithms for $k$-HodgeRank. Known block encodings of $\bm{\mathrm{B}}_{k}$ are constructed via quantum random access memory (QRAM) or a sparse access input model (SAIM). In these cases, $\alpha_{\bm{\mathrm{B}}_{k}}$ scales with either the Frobenius norm or the sparsity of $\bm{\mathrm{B}}_{k}$ [37, 39], which both have worst-case scaling $\Omega(n^{k})$ as discussed previously. We summarize these relationships between classical least squares algorithms, quantum least squares algorithms, HodgeRank, and quantum $k$-HodgeRank in Figure 1 in Appendix III.
@@
where $\lvert\sigma_{k}^{i}\rangle$ is an $n$-qubit computational basis state representing $\sigma_{k}^{i}$ defined in [45, 43] (given in Eq. (18) in Appendix II). We discuss the complexities of our applications in terms of the number of calls to $\bm{\mathrm{U}}_{\mathrm{prep}}$.
QTSP combines the quantum singular value transformation (QSVT) [41, 46] and PUEs $\bm{\mathrm{U}}_{\bm{\mathrm{\bm{\mathrm{B}}_{k}}}}$ and $\bm{\mathrm{U}}_{\bm{\mathrm{\bm{\mathrm{B}}_{k+1}}}}$ of the boundary matrices $\bm{\mathrm{B}}_{k}$ and $\bm{\mathrm{B}}_{k+1}$ [42, 47] (see Appendix II). It prepares, upon successful postselection of the ancilla (which occurs with probability approximately $\mathcal{N}^{2}:=\big{\|}H\left(\bm{\mathrm{B}}_{k}/\sqrt{n},\bm{\mathrm{B}}_{k+1}^{\dagger}/\sqrt{n}\right)\lvert\bm{\mathrm{s}}^{k}\rangle\big{\|}_{2}^{2}$), the filtered state
@@
Here $H(x,y):[-1,1]^{2}\rightarrow[-1,1]$ is a sum of two real polynomials $p(x)$ and $q(y)$, each with degree at most $D$, applied to the singular values of the respective matrices, where $p$ and $q$ are either both even or one is zero. The polynomial $H(x,y)$ is chosen so that |$\ket{\bm{\mathrm{s}}^{k}_{\mathrm{fil}}}$ extracts specific information from |$\ket{\bm{\mathrm{s}}^{k}}$. QTSP requires one call to $\bm{\mathrm{U}}_{\mathrm{prep}}$ as well as $O\big{(}D\big{)}$ calls to $\bm{\mathrm{U}}_{\bm{\mathrm{B}}_{k}}$, $\bm{\mathrm{U}}_{\bm{\mathrm{B}}_{k+1}}$, their inverses, and $O(D)$ applications of $\mathrm{C}_{\Pi_{j}}\mathrm{NOT}$ for $j\in\{k-1,k,k+1\}$, where $\Pi_{j}$ is the projector onto the states corresponding to $j$-simplices [48, 49], as well as post-selection on the correct ancilla state. The $\mathrm{C}_{\Pi_{j}}\mathrm{NOT}$ gates depend on the edges in the underlying graph $\mathcal{G}$ of the clique complex $\mathcal{K}_{n}(\mathcal{G})$ and are implemented via classical control. The circuit for QTSP has non-Clifford gate depth $O(Dn\log n)$ and uses $O(n)$ qubits. We review QTSP and its complexity in Appendix II (see [44]).
@@
This is a $2\varepsilon/\left(\mathcal{N}_{\ast}-\varepsilon\right)$-approximation to $\lvert\bm{\mathrm{s}}^{k-1}_{\ast}\rangle$ (in $\ell_{2}$ norm) if $\mathcal{N}_{*}>\varepsilon$. In cases where $\bm{\mathrm{s}}^{k}$ has high rankability (defined formally in Application 1), $\varepsilon$ can be chosen such that $\|\lvert\tilde{\bm{\mathrm{s}}}^{k-1}_{\ast}\rangle-\lvert{\bm{\mathrm{s}}}^{k-1}_{\ast}\rangle\|\leq 4\varepsilon/\mathcal{N}_{\ast}$ without significantly affecting the complexity (see Appendix III). The complexity of QTSP implies that preparing |$\ket{\tilde{\bm{\mathrm{s}}}^{k-1}_{*}}$ from |$\ket{\bm{\mathrm{s}}^{k}}$, excluding post-selection, uses $O(n)$ qubits, a single call to $\bm{\mathrm{U}}_{\mathrm{prep}}$, and has non-Clifford gate depth
@@
If $\kappa_{k}$ is small (e.g. not growing exponentially in $k$), this is much more efficient than classically computing $\bm{\mathrm{s}}^{k-1}_{*}$, even for small $k$. Examples of simplicial complexes with slowly-vanishing (or even constant in $k$) spectral gap are given in [43] and discussed in Appendix VII. Computing $p(\bm{\mathrm{B}}_{k})\bm{\mathrm{s}}^{k}$ classically requires $\tilde{O}(\kappa_{k}^{2})$ iterations of matrix-vector multiplication that each require $O(n_{k})$ operations [36]. This again highlights the dependency of the simplex dimension in the classical complexity, particularly for complexes with many $k$-simplices, i.e. when $n_{k}=\Theta\left(\binom{n}{k+1}\right)\approx\Theta(n^{k+1})$.
@@
The value $\mathrm{R}(k)$ quantifies whether (sets of) alternatives can be meaningfully ranked based on the data, where larger values indicate high rankability. Conversely, $\mathrm{R}_{\mathrm{C}}(k)$ measures the presence of non-transitive preference relationships in the data, where the consensus prefers $A$ over $B$, $B$ over $C$, but $C$ over $A$. This has applications in finding arbitrage in currency exchange markets [4]. These measures are related to what is known as the Hodge decomposition (see Appendix I for a brief introduction to the underlying topology).
We provide algorithms to compute $\varepsilon$-approximations to $\mathrm{R}(k)$ and $\mathrm{R}_{\mathrm{C}}(k)$ (see Appendix IV for details). Eq. (9) implies that $\mathrm{R}(k)=\left(\bra{\bm{\mathrm{s}}^{k}}\Pi_{\mathrm{G}}\ket{\bm{\mathrm{s}}^{k}}\right)^{1/2}$, where $\Pi_{\mathrm{G}}:=\bm{\mathrm{B}}_{k}^{\dagger}(\bm{\mathrm{B}}_{k}\bm{\mathrm{B}}_{k}^{\dagger})^{+}\bm{\mathrm{B}}_{k}$ projects onto $\mathrm{im}(\bm{\mathrm{B}}_{k}^{\dagger})$. Thus, we approximate $\mathrm{R}(k)$ using an accelerated Hadamard test [52] by using QTSP to build a PUE of (an approximate) $\Pi_{\mathrm{G}}$. Let $H(x,y)=p(x)=x^{2}g_{\varepsilon^{2}/2}(x^{2})$, where as before $g_{\varepsilon}$ is a polynomial which $\varepsilon/(2\kappa_{k}^{2})$-approximates $1/(2\kappa_{k}^{2}x)$. Then $\big{\|}\Pi_{\mathrm{G}}-2\kappa_{k}^{2}\,p\big{(}\bm{\mathrm{B}}_{k}/\sqrt{n}\big{)}\big{\|}_{2}\leq\varepsilon^{2}/2$, where $p(x)$ is degree $D=O\big{(}\kappa_{k}^{2}\log(n\kappa_{k}^{2}/\varepsilon^{2})\big{)}$ (see [44], Theorem 2). Thus, if $\bm{\mathrm{U}}_{\Pi_{\mathrm{G}}}=\mathrm{QTSP}(k,\mathcal{K}_{n},x^{2}g_{\varepsilon^{2}/2}(x^{2}))$ is the unitary that block encodes $p\big{(}\bm{\mathrm{B}}_{k}/\sqrt{n}\big{)}$ in the subspace defined by the ancilla in state |$\ket{0}^{\otimes a}$, then
@@
Relative rankings of few alternatives.—We can use quantum $k$-HodgeRank to approximate the scores of individual alternatives. This gives a heuristic for relatively ranking a small subset of alternatives without full-state tomography or computing classical $k$-HodgeRank. For more details see Appendix V.
@@
Finding good alternatives.— We now discuss determining an approximate score for every $(k-1)$-simplex. As mentioned previously, full-state tomography is expensive. However, the tomography techniques given in [53] returns a classically stored $\ell_{\infty}$-norm approximation to |$\ket{\tilde{\bm{\mathrm{s}}}^{k-1}_{*}}$, denoted $\hat{\bm{\mathrm{s}}}^{k-1}_{*}$. As detailed in Appendix VI, one can determine the score of every $(k-1)$-simplex, up to additive error $O(\varepsilon)$, in $O(\kappa_{k}^{4}n\varepsilon^{-2})$ applications of $\mathrm{QTSP}(k,\mathcal{K}_{n},xg_{\varepsilon}(x))$ and $\bm{\mathrm{U}}_{\mathrm{prep}}$. This $\ell_{\infty}$-norm approximation can be used to find alternatives with scores within additive error $O(\varepsilon)$ of the best alternative, or to determine whether the alternatives can be classified into subsets with similar scores within subsets and large differences in scores between subsets. However, like the previous application this has issues as the amplitudes of the state |$\ket{\bm{\mathrm{s}}^{k-1}_{*}}$ may be arbitrarily close, and an appropriate choice of $\varepsilon$ cannot be easily determined from |$\ket{\bm{\mathrm{s}}^{k}}$.
Dequantization. — Chia et al. [54] and Gharibian and Le Gall [55] have given different frameworks for dequantizing algorithms based on QSVT. In Appendix VII we show that these frameworks do not generically lead to efficient classical dequantizations of quantum $k$-HodgeRank or the computation of consistency or local inconsistency scores (Application 1). The reasons for this are: (i) difficulties obtaining polylogarithmic-time sampling and query access to $\bm{\mathrm{B}}_{k}$, and (ii) there exist simplicial complexes such that $\bm{\mathrm{B}}_{k}$ has Frobenius norm $\Theta(n^{k})$ and row sparsity $\Theta(n)$. We present a family of such complexes, previously studied by Berry et al. [43] in the context of quantum topological data analysis (QTDA). Each of these issues leads to inefficient dequantized algorithms in these frameworks. Methods for dequantizing QTDA rely on the relationship to estimating the dimension of the null space of an operator and do not obviously apply to our algorithms. An interesting future direction would be determining to what extent QTSP and quantum $k$-HodgeRank can be dequantized.

### v3

@@ Abstract
where $w_{ij}$ represents the weight associated to the data on (oriented) edge $[v_{i},v_{j}]$ and $\bm{\mathrm{B}}_{k}$ is a matrix representation of the $k$-th boundary operator of a clique complex $\mathcal{K}_{n}(\mathcal{G})$ (see Supplemental Material [32]). The space $\mathrm{im}(\bm{\mathrm{B}}_{1}^{\dagger})$ contains so-called conservative edge flows, which means there exists $\bm{\mathrm{s}}^{0}_{*}\in\mathbb{R}^{|\mathcal{V}|}$ such that $\bm{\mathrm{s}}^{1}_{\mathrm{G}}([v_{i},v_{j}])=\bm{\mathrm{s}}^{0}_{*}(v_{i})-\bm{\mathrm{s}}^{0}_{*}(v_{j})$. In a discrete exterior calculus sense, $\bm{\mathrm{s}}^{1}_{\mathrm{G}}$ is an analogue of the gradient of the potential function $\bm{\mathrm{s}}^{0}_{*}$.
HodgeRank then outputs the potential (or score) function $\bm{\mathrm{s}}^{0}_{*}$ from Eq. (1) by $\bm{\mathrm{B}}_{1}^{\dagger}\bm{\mathrm{s}}^{0}_{\ast}=\bm{\mathrm{s}}^{1}_{\mathrm{G}}$. This defines a global ranking of the alternatives: the score of alternative $v_{i}$ is $\bm{\mathrm{s}}^{0}_{*}(v_{i})$, and $v_{i}$ outranks $v_{j}$ whenever $\bm{\mathrm{s}}^{0}_{*}(v_{i})\geq\bm{\mathrm{s}}^{0}_{*}(v_{j})$. In this Letter, we set $w_{ij}=0$ if $v_{i},v_{j}$ are not compared, and 1 otherwise; we discuss this assumption in Supplemental Material [32]. Under this assumption, $\bm{\mathrm{s}}^{0}_{\ast}=\big{(}\bm{\mathrm{B}}_{1}\bm{\mathrm{B}}_{1}^{\dagger}\big{)}^{+}\bm{\mathrm{B}}_{1}\bm{\mathrm{s}}^{1}$, where $+$ denotes the Moore-Penrose pseudoinverse. The relative length of $\bm{\mathrm{s}}^{1}_{\mathrm{G}}$ compared to $\bm{\mathrm{s}}^{1}$ measure the rankability of the alternatives according to $\bm{\mathrm{s}}^{1}$.
@@
These algorithms do not lead to efficient quantum algorithms for $k$-HodgeRank. Known block encodings of $\bm{\mathrm{B}}_{k}$ are constructed via quantum random access memory (QRAM) or a sparse access input model (SAIM). In these cases, $\alpha_{\bm{\mathrm{B}}_{k}}$ scales with either the Frobenius norm or the sparsity of $\bm{\mathrm{B}}_{k}$ [38, 40], which both have worst-case scaling $\Omega(n^{k})$ as discussed previously. We summarize these relationships between classical least squares algorithms, quantum least squares algorithms, HodgeRank, and quantum $k$-HodgeRank in Figure 1 in Supplemental Material [32].
@@
where $\lvert\sigma_{k}^{i}\rangle$ is an $n$-qubit computational basis state representing $\sigma_{k}^{i}$ defined in [46, 44] (given in Eq. (18) in Supplemental Material [32]). We discuss the complexities of our applications in terms of the number of calls to $\bm{\mathrm{U}}_{\mathrm{prep}}$.
QTSP combines the quantum singular value transformation (QSVT) [42, 47] and PUEs $\bm{\mathrm{U}}_{\bm{\mathrm{\bm{\mathrm{B}}_{k}}}}$ and $\bm{\mathrm{U}}_{\bm{\mathrm{\bm{\mathrm{B}}_{k+1}}}}$ of the boundary matrices $\bm{\mathrm{B}}_{k}$ and $\bm{\mathrm{B}}_{k+1}$ [43, 48] (see Supplemental Material [32]). It prepares, upon successful postselection of the ancilla (which occurs with probability approximately $\mathcal{N}^{2}:=\big{\|}H\left(\bm{\mathrm{B}}_{k}/\sqrt{n},\bm{\mathrm{B}}_{k+1}^{\dagger}/\sqrt{n}\right)\lvert\bm{\mathrm{s}}^{k}\rangle\big{\|}_{2}^{2}$), the filtered state
@@
Here $H(x,y):[-1,1]^{2}\rightarrow[-1,1]$ is a sum of two real polynomials $p(x)$ and $q(y)$, each with degree at most $D$, applied to the singular values of the respective matrices, where $p$ and $q$ are either both even or one is zero. The polynomial $H(x,y)$ is chosen so that |$\ket{\bm{\mathrm{s}}^{k}_{\mathrm{fil}}}$ extracts specific information from |$\ket{\bm{\mathrm{s}}^{k}}$. QTSP requires one call to $\bm{\mathrm{U}}_{\mathrm{prep}}$ as well as $O\big{(}D\big{)}$ calls to $\bm{\mathrm{U}}_{\bm{\mathrm{B}}_{k}}$, $\bm{\mathrm{U}}_{\bm{\mathrm{B}}_{k+1}}$, their inverses, and $O(D)$ applications of $\mathrm{C}_{\Pi_{j}}\mathrm{NOT}$ for $j\in\{k-1,k,k+1\}$, where $\Pi_{j}$ is the projector onto the states corresponding to $j$-simplices [49, 50], as well as post-selection on the correct ancilla state. The $\mathrm{C}_{\Pi_{j}}\mathrm{NOT}$ gates depend on the edges in the underlying graph $\mathcal{G}$ of the clique complex $\mathcal{K}_{n}(\mathcal{G})$ and are implemented via classical control. The circuit for QTSP has non-Clifford gate depth $O(Dn\log n)$ and uses $O(n)$ qubits. We review QTSP and its complexity in Supplemental Material [32] (see [45]).
@@
This is a $2\varepsilon/\left(\mathcal{N}_{\ast}-\varepsilon\right)$-approximation to $\lvert\bm{\mathrm{s}}^{k-1}_{\ast}\rangle$ (in $\ell_{2}$ norm) if $\mathcal{N}_{*}>\varepsilon$. In cases where $\bm{\mathrm{s}}^{k}$ has high rankability (defined formally in Application 1), $\varepsilon$ can be chosen such that $\|\lvert\tilde{\bm{\mathrm{s}}}^{k-1}_{\ast}\rangle-\lvert{\bm{\mathrm{s}}}^{k-1}_{\ast}\rangle\|\leq 4\varepsilon/\mathcal{N}_{\ast}$ without significantly affecting the complexity (see Supplemental Material [32]). The complexity of QTSP implies that preparing |$\ket{\tilde{\bm{\mathrm{s}}}^{k-1}_{*}}$ from |$\ket{\bm{\mathrm{s}}^{k}}$, excluding post-selection, uses $O(n)$ qubits, a single call to $\bm{\mathrm{U}}_{\mathrm{prep}}$, and has non-Clifford gate depth
@@
If $\kappa_{k}$ is small (e.g. not growing exponentially in $k$), this is much more efficient than classically computing $\bm{\mathrm{s}}^{k-1}_{*}$, even for small $k$. Examples of simplicial complexes with slowly-vanishing (or even constant in $k$) spectral gap are given in [44] and discussed in Supplemental Material [32]. Computing $p(\bm{\mathrm{B}}_{k})\bm{\mathrm{s}}^{k}$ classically requires $\tilde{O}(\kappa_{k}^{2})$ iterations of matrix-vector multiplication that each require $O(n_{k})$ operations [37]. This again highlights the dependency of the simplex dimension in the classical complexity, particularly for complexes with many $k$-simplices, i.e. when $n_{k}=\Theta\left(\binom{n}{k+1}\right)\approx\Theta(n^{k+1})$.
@@
The value $\mathrm{R}(k)$ quantifies whether (sets of) alternatives can be meaningfully ranked based on the data, where larger values indicate high rankability. Conversely, $\mathrm{R}_{\mathrm{C}}(k)$ measures the presence of non-transitive preference relationships in the data, where the consensus prefers $A$ over $B$, $B$ over $C$, but $C$ over $A$. This has applications in finding arbitrage in currency exchange markets [4]. These measures are related to what is known as the Hodge decomposition (see Supplemental Material [32] for a brief introduction to the underlying topology).
We provide algorithms to compute $\varepsilon$-approximations to $\mathrm{R}(k)$ and $\mathrm{R}_{\mathrm{C}}(k)$ (see Supplemental Material [32] for details). Eq. (9) implies that $\mathrm{R}(k)=\left(\bra{\bm{\mathrm{s}}^{k}}\Pi_{\mathrm{G}}\ket{\bm{\mathrm{s}}^{k}}\right)^{1/2}$, where $\Pi_{\mathrm{G}}:=\bm{\mathrm{B}}_{k}^{\dagger}(\bm{\mathrm{B}}_{k}\bm{\mathrm{B}}_{k}^{\dagger})^{+}\bm{\mathrm{B}}_{k}$ projects onto $\mathrm{im}(\bm{\mathrm{B}}_{k}^{\dagger})$. Thus, we approximate $\mathrm{R}(k)$ using an accelerated Hadamard test [53] by using QTSP to build a PUE of (an approximate) $\Pi_{\mathrm{G}}$. Let $H(x,y)=p(x)=x^{2}g_{\varepsilon^{2}/2}(x^{2})$, where as before $g_{\varepsilon}$ is a polynomial which $\varepsilon/(2\kappa_{k}^{2})$-approximates $1/(2\kappa_{k}^{2}x)$. Then $\big{\|}\Pi_{\mathrm{G}}-2\kappa_{k}^{2}\,p\big{(}\bm{\mathrm{B}}_{k}/\sqrt{n}\big{)}\big{\|}_{2}\leq\varepsilon^{2}/2$, where $p(x)$ is degree $D=O\big{(}\kappa_{k}^{2}\log(n\kappa_{k}^{2}/\varepsilon^{2})\big{)}$ (see [45], Theorem 2). Thus, if $\bm{\mathrm{U}}_{\Pi_{\mathrm{G}}}=\mathrm{QTSP}(k,\mathcal{K}_{n},x^{2}g_{\varepsilon^{2}/2}(x^{2}))$ is the unitary that block encodes $p\big{(}\bm{\mathrm{B}}_{k}/\sqrt{n}\big{)}$ in the subspace defined by the ancilla in state |$\ket{0}^{\otimes a}$, then
@@
Relative rankings of few alternatives.—We can use quantum $k$-HodgeRank to approximate the scores of individual alternatives. This gives a heuristic for relatively ranking a small subset of alternatives without full-state tomography or computing classical $k$-HodgeRank. For more details see Supplemental Material [32].
@@
Finding good alternatives.— We now discuss determining an approximate score for every $(k-1)$-simplex. As mentioned previously, full-state tomography is expensive. However, the tomography techniques given in [54] returns a classically stored $\ell_{\infty}$-norm approximation to |$\ket{\tilde{\bm{\mathrm{s}}}^{k-1}_{*}}$, denoted $\hat{\bm{\mathrm{s}}}^{k-1}_{*}$. As detailed in Supplemental Material [32], one can determine the score of every $(k-1)$-simplex, up to additive error $O(\varepsilon)$, in $O(\kappa_{k}^{4}n\varepsilon^{-2})$ applications of $\mathrm{QTSP}(k,\mathcal{K}_{n},xg_{\varepsilon}(x))$ and $\bm{\mathrm{U}}_{\mathrm{prep}}$. This $\ell_{\infty}$-norm approximation can be used to find alternatives with scores within additive error $O(\varepsilon)$ of the best alternative, or to determine whether the alternatives can be classified into subsets with similar scores within subsets and large differences in scores between subsets. However, like the previous application this has issues as the amplitudes of the state |$\ket{\bm{\mathrm{s}}^{k-1}_{*}}$ may be arbitrarily close, and an appropriate choice of $\varepsilon$ cannot be easily determined from |$\ket{\bm{\mathrm{s}}^{k}}$.
Dequantization. — Chia et al. [55] and Gharibian and Le Gall [56] have given different frameworks for dequantizing algorithms based on QSVT. In Supplemental Material [32] we show that these frameworks do not generically lead to efficient classical dequantizations of quantum $k$-HodgeRank or the computation of consistency or local inconsistency scores (Application 1). The reasons for this are: (i) difficulties obtaining polylogarithmic-time sampling and query access to $\bm{\mathrm{B}}_{k}$, and (ii) there exist simplicial complexes such that $\bm{\mathrm{B}}_{k}$ has Frobenius norm $\Theta(n^{k})$ and row sparsity $\Theta(n)$. We present a family of such complexes, previously studied by Berry et al. [44] in the context of quantum topological data analysis (QTDA). Each of these issues leads to inefficient dequantized algorithms in these frameworks. Methods for dequantizing QTDA rely on the relationship to estimating the dimension of the null space of an operator and do not obviously apply to our algorithms. An interesting future direction would be determining to what extent QTSP and quantum $k$-HodgeRank can be dequantized.